An acidophilus culture containing 150 bacteria doubles every 20 minutes. Write the exponential function then determine how much bacteria there are after 4 hours.


1
Expert's answer
2021-03-09T06:03:42-0500

"20 \ min=\dfrac{1}{3}\ h"

Function that represents the number of bacteria is


"N(t)=N(0)(2^{3t})"

Given "N(0)=150, t=4\ h"


"N(4)=150(2^{3(4)})=614400"

There are 614400 bacteria after 4 hours